. Consider a shop that sells mobile phones. The number of customers arriving per
hour is assumed to follow a Poisson distribution with mean 2. Each customer
visiting the shop will buy a phone with probability 0.25. Answer the following
questions:
(a) What is the probability that more than 3 customers will arrive during the next
hour?
(b) What is the probability that in a 2-hour period not more than 5 customers
will arrive?
2
(c) If 15 customers visit the shop during a day, what is the probability that they
will sell at least 5 phones?
(d) When the shop opens in the morning, what is the probability that they have
to wait more than half an hour for the first customer to arrive?
1
Expert's answer
2019-11-25T11:55:20-0500

a) "\lambda=2."

"P(X>3)=1-P(X=0)-P(X=1)-P(X=2)-P(X=3)=\\
=1-e^{-2}(\frac{2^0}{0!}+\frac{2^1}{1!}+\frac{2^2}{2!}+\frac{2^3}{3!})=0.1429."


b) "\lambda=2*2=4."

"P(X>5)=1−P(X=0)−P(X=1)−P(X=2)−P(X=3)+\\+P(X=4)+P(x=5)=\\
=1−e^{−4}(\frac{4^0}{0!}+\frac{4^1}{1!}+\frac{4^2}{2!}+\frac{4^3}{3!}+
\frac{4^4}{4!}+\frac{4^5}{5!})=0.2149."


c) Binomial distribution with "n=15, p=0.25."

"P(X\ge5)=1-P(X<5)=1-C_{15}^00.25^00.75^5-C_{15}^10.25^10.75^4-\\
-C_{15}^20.25^20.75^3-C_{15}^30.25^30.75^2-C_{15}^40.25^40.75^1=0.3135."


d) There are no customers in the first half of hour "\lambda=\frac{2}{2}=1" .

"P(X=0)=e^{-1}\frac{1^0}{0!}=0.3679."